OpenStudy (anonymous):

Estimation Estimate each sum or difference 25 5/18 - 9 11/17

OpenStudy (anonymous):



OpenStudy (anonymous):

Well, do we need to solve this and then estimate or are these two different numbers and we just need to estimate the two different numbers? Ok, to estimate these two numbers we can take the fraction and then turn it into a decimal. For instance: 25 5/18 5/18 = .277 25 + .277 = 25.277 25.277 = 25 if we are estimating to the nearest ones. 25.277 = 25.20 if we are estimating to the nearest tenths 25.277 = 25.28 if we are estimating to the nearest hundredths Can you do the next one?

OpenStudy (anonymous):

i dont really understand

OpenStudy (anonymous):

Does your direction tell you what to estimate to? Are you to just estimate to the nearest ones?

OpenStudy (anonymous):

Do you know how to take a fraction and turn it into a decimal?

OpenStudy (anonymous):

it just say estimate each sum or difference

OpenStudy (anonymous):

Ok, does it give you answers to chose from?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

For -9 11/17 -9 11/17 11/17 = .647 -9 + .647 = -9.647 -9.647 = -10 if we are estimating to the nearest ones. -9.647 = -9.60 if we are estimating to the nearest tenths -9.647 = -9.65 if we are estimating to the nearest hundredths You pretty much are turning the fraction into a decimal and then rounding to what ever place they want you to round / estimate to. I am assuming the ones place so your answer would be 25 and -10 for the two numbers

OpenStudy (anonymous):

how do i get answer

OpenStudy (anonymous):

You need to take the fraction and turn it into a deciaml. Once you have the decimal you need to round the number. Do you know how to round numbers and turn a fraction into decimals?

OpenStudy (anonymous):

Rounding numbers http://www.mathsisfun.com/rounding-numbers.html Fractions to decimals http://www.mathsisfun.com/converting-fractions-decimals.html

OpenStudy (anonymous):

You all might be doing it like this in school. https://www.youtube.com/watch?v=TYgNtYr8n8I this is just estimating the fraction with out turning the fraction into a decimal